Obtain the volume of rectangular boxes with the following length, breadth and height respectively
xy, 2x^2y, 2xy^2

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Volume of rectangle = length x breadth x height.
To evaluate volume of rectangular boxes, multiply all the monomials.
xy, 2x^2y, 2xy^2= (1×2× 2)( x×x^2×x×y×y×y^2 ) =4x^4y^4

​Thanks